OpenStudy (ammarah):

Write an equation of the parabola with the vertex at (5,-3) and focus at (1,-3) @ganeshie8

OpenStudy (ammarah):

(y-k)^2=4p(x-h)

ganeshie8 (ganeshie8):

First, figure out whether it faces UP-DOWN, or LEFT-RIGHT

OpenStudy (ammarah):

left and right?

OpenStudy (ammarah):

cuz the foci

ganeshie8 (ganeshie8):

Correct ! its a left and right since the y coordinate is fixed

ganeshie8 (ganeshie8):

and you know vertex already, so the equation will look like : \(\large (y-(-3))^2=4p(x-5)\) right ?

ganeshie8 (ganeshie8):

you just need to find the value of \(\large p\)

OpenStudy (ammarah):

yes

OpenStudy (ammarah):

wait so is that it for the equation cuz its only asking for the equation

ganeshie8 (ganeshie8):

you still need to find the value of \(\large p\) p = focus - vertex = 1 - 5 = -4

ganeshie8 (ganeshie8):

so, the equation wud be : \(\large (y-(-3))^2=4(-4)(x-5) \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (ammarah):

ok

OpenStudy (ammarah):

would it be y+3? or do i just leave it like that

ganeshie8 (ganeshie8):

either way is fine...

ganeshie8 (ganeshie8):

\(\large (y-(-3))^2=4(-4)(x-5)\) is same as : \(\large (y+3)^2=-16(x-5)\)